OpenStudy (anonymous):

wich of the numbers listed below are solutions to the equation x2=49 a. -7 b.2401 c.24.5 d.7 e.14 f.-14 ?? so confused help pla can some one explain this one ?

OpenStudy (anonymous):

replace x with each one of those and see if you get a true statement....

OpenStudy (anonymous):

\[x^2=49\] square both sides \[\sqrt{x^2}=\sqrt{49}\] \[x=7\]

OpenStudy (anonymous):

sorry \[x=\pm7\]